OpenStudy (anonymous):

Simplify, Assume all variables are positive. squrt245p^2q^5r I need help with this problem I don't really get the processes can any one do this out please

OpenStudy (anonymous):

use the equation button below to type out your question clearly. | | | v

OpenStudy (anonymous):

\[\sqrt{245p^2q^5r}\]

OpenStudy (anonymous):

\[=\sqrt{7^2\times 5\times p^2\times q^5\times r}\] \[=\sqrt{7^2}\times \sqrt{p^2}\times \sqrt{q^4}\times \sqrt{5qr}\] \[=7pq^2\sqrt{5qr}\]

OpenStudy (anonymous):

\[\sqrt{252p^2q^9r}\]

OpenStudy (anonymous):

right now how did you get those answers.

OpenStudy (anonymous):

by the definition of square root

OpenStudy (anonymous):

which is?

OpenStudy (anonymous):

\[\sqrt{p^2} = p\]

OpenStudy (anonymous):

gimmick is to find all the perfect squares inside the radical, and pull them out. so if you see \[\sqrt{50}\] you think \[\sqrt{2\times 25}=5\sqrt{2}\]

OpenStudy (anonymous):

and if you see \[\sqrt{x^5}\] you think \[\sqrt{x^4\times x}=x^2\sqrt{x}\]

OpenStudy (anonymous):

Perfect thanks!